# Symmetry operations and Critical Behaviour in Classical to Quantum Stochastic Processes

### v1

@@ Abstract
Construction of quantum analogs starting from classical stochastic processes have been previously introduced. In this paper, we generate a large class of self contained quantum extensions by symmetry operations. We show that the relaxation processes for different quantum extensions are different and that is supported by the measure of coherence, the the probability of reaching the equilibrium, decay of the domain walls and purity. However, the coherence measure based on the L1-norm does not capture the speed of the relaxation process. We also show that the finite size scaling of coherence exists for both short and long times.
@@ I Introduction
Here we investigate the frontier between quantum systems and classical stochastic systems. On a fundamental level, stochasticity is not intrinsic to classical dynamics, it rather emerges due to the chaotic dynamics of many particle systems. In quantum mechanics, by contrast, genuine stochasticity appears as soon as measurement processes are included (wave function collapse). Even though the measurement process can be avoided in many different ways, for instance by including the measurement apparatus into the quantum system to be described, at some point, it has to pop up, if meaningful experimental results are to be discussed.
@@
A gradual transition from quantum dynamics (deterministic and unitary) to (partially) stochastic processes is usually the subject of study in the theory of open quantum systems. In this case, information about the dynamics of the quantum system is leaking out into the environment, a process which may be seen as applying measurements continuously with low probabilities.
The characterization of such Markovian processes has been established in Refs. [1, 2, 3], and is by now well established and used in countless applications.
@@
Recently, an alternative approach (to describe systems intermediate between quantum unitary and classical stochastic) has been proposed [4]. There, we start from the classical stochastic process, and ask which are the (partially) quantum processes in its vicinity (quantum extensions). Surprisingly, this approach lead to processes which are close to classical (but still quantum, with low coherence – measured by the L1-norm based measure introduced in [5]). However, it shows macroscopic behavior that is different (even in the scaling exponents) as compared to the classical case.
@@
The finding of stochastic processes which are close to classical (in terms of the coherence measure) and still show very different properties on the macroscopic scale (thermodynamic limit) is important for many reasons: (i) bio-chemical processes achieving .. efficiencies (light harvesting, etc.) [6, 7] (ii) Superior quantum computation with low-coherence quantum computers, to name two [8].
@@
While this first paper was meant as a proof of principle, in the present contribution, we investigate quantum extensions in a broader context. Hence, we construct a large self-contained set of quantum extensions, based on a single (two) symmetry operations. For the relaxation process in the Ising model, we investigate the speed of the process by measuring several quantities as a function of time: (i) the probability to find the system in the ground state; (ii) the decay of the number of domains. We study the behavior of the coherence with time and show the finite size scaling for its decay exist. We find that the amount of coherence is related to the degree to which the quantum extension deviates from the purely classical process. Finally, we study the evolution of the purity of the system.
@@ II Classical and quantum stochastic processes
Here we consider classical stochastic processes with a finite and discrete sample space, $\{|j\rangle\}_{j=1,\ldots,d}$, where each $|j\rangle$ represents a possible configuration of the physical system. In the case of a discrete sample space, the dynamics of the system can be characterized by a sequence of stochastic transition matrices, $\{\mathcal{T}(n)\,|\,n\in\mathbb{N}\}$, such that
$\bm{p}(n+1)=\mathcal{T}(n)\;\bm{p}(n)\quad:\quad p_{i}(m+1)=\sum_{j}\mathcal{T}_{ij}(m)\;p_{j}\;.$ (1)
Where the $j$th element, $p_{j}(n)$, of the vector $\bm{p}$ denotes the probability that the system is found in the configuration $|j\rangle$ at the discrete time $n$. Each element $\mathcal{T}_{ij}\geq 0$, of the matrix $\mathcal{T}$ of dimensions $d\times d$, is a conditional probability such that each $j$th column of the transition matrix satisfy the condition, $\sum_{i}\mathcal{T}_{ij}=1$. This guarantees that for every discrete time $n$ the total probability is conserved, i.e. $\sum_{j}p_{j}(n)=1$. The stochastic process may then be understood as the following composition of transition matrices
$\bm{p}(n)=[\mathcal{T}(n)\circ\mathcal{T}(n-1)\circ\cdots\circ\mathcal{T}(m)]\bm{p}(m-1)\,.$ (2)
Where the composition law, $[\mathcal{T}(n)\circ\mathcal{T}(n-1)\circ\cdots\circ\mathcal{T}(m)]_{ij}$, express the probability to find the system in state $|i\rangle$ at time $n$, provided the system was in the state $|j\rangle$ at time $m-1$ [9].
@@ II.1 Quantum processes
From the point of view of open quantum systems, the ensembles are denoted by density matrices $\varrho\in\mathcal{S}(\mathcal{H})$ [10, 11]. Here, $\mathcal{H}$ is the Hilbert space defined as the space of all linear combinations of classical states $\{|j\rangle\}_{j=1,\ldots,d}$, with a scalar product satisfying the condition $\langle i|j\rangle=\delta_{ij}$. $\mathcal{S}(\mathcal{H})$ stands for the space of all convex combination of density matrices. A Markovian quantum stochastic process is defined as a sequence of quantum maps, $\Lambda_{\mathcal{Q}}$, which satisfy the condition of Complete Positivity and Trace Preserving (CPTP), in the space $\mathcal{S}$ [12]. Now, instead of evolving probability vectors, density matrices are evolved in time, according to the following composition law [13, 14],
$\varrho(n)$ $=\Lambda_{\mathcal{Q}}(n)\circ\Lambda_{\mathcal{Q}}(n-1)\circ\cdots\circ\Lambda_{\mathcal{Q}}(m)[\varrho(m-1)]$ (3)
where $\varrho(m-1)$ and $\varrho(n)$ stands for the initial and final ensembles respectively. Considering that every CPTP map admits a sum representation of Kraus operators [15]. Each element in the sequence, Eq.(3), is written as follows
$\Lambda_{\mathcal{Q}}(n)[\varrho(n-1)]$ $=\sum_{\alpha}\mathcal{K}_{\alpha}(n)\,\varrho(n-1)\,\mathcal{K}_{\alpha}^{\dagger}(n)\,,$ (4)
where the Kraus operators, $\{\mathcal{K}_{\alpha}\}$, satisfy the condition
$\sum_{\alpha}\mathcal{K}_{\alpha}^{\dagger}\mathcal{K}_{\alpha}=\mathds{1}.$ (5)
Interestingly, even classical stochastic processes, as those in Eq.(2) can be represented in this formalism of open quantum systems. First, by expressing the probability vector, $\bm{p}$, as a diagonal density matrix, $\bar{\varrho}$. Subsequently, by replacing the transition matrix, $\mathcal{T}$, by a map $\Lambda_{\mathcal{T}}$,
$\bar{\varrho}(n)=\Lambda_{\mathcal{T}}(n)[\bar{\varrho}(n-1)]=\sum_{i,j}K_{ij}(n)\bar{\varrho}(n-1)K_{ij}^{\dagger}(n),$ (6)
with the following structure for the Kraus operators
$\Big{\{}K_{ij}(n)$ $=\sqrt{\mathcal{T}_{ij}(n)}{|i\rangle\langle j|}\Big{\}}_{ij}\,.$ (7)
From the above definition we know two things, firstly that condition (5) is always satisfied, and therefore, $\Lambda_{\mathcal{T}}$ is a valid CPTP map; secondly, that the evolution defined in Eq.(6), is purely classical. Since the map $\Lambda_{\mathcal{T}}$, belongs to the set of incoherent CPTP maps, i.e. , which are neither able to create nor detect coherence [5, 16].
@@ II.2 Quantum extensions
We are interested in quantum processes that can be reduced to classical processes when they are observed or measured sufficiently often. We name such processes as quantum extensions of the classical stochastic process. This idea can be clarified by the following formal definition
Definition. [4]
@@
where $\mathcal{P}$ denotes a complete measurement of the set of classical states (i.e. , the basis states $|j\rangle$).
To find quantum processes, such as the one defined in Eq.(3), that can actually be reduced to a classical stochastic process, such as Eq.(2), or equivalently
$\bar{\varrho}(n)$ $=\Lambda_{\mathcal{T}}(n)\circ\Lambda_{\mathcal{T}}(n-1)\circ\cdots\circ\Lambda_{\mathcal{T}}(m)[\bar{\varrho}(m-1)]$ (9)
we use the following guiding principles. (i) We try to find quantum extensions for each stochastic map $\Lambda_{\mathcal{T}}(n)$, individually. (ii) Both the classical and the quantum map can be uniquely defined by their action on the basis states $\{|j\rangle\}$. Here, the classical map is determined by the probabilities to jump from the state $|j\rangle$ to any other state $|i\rangle$. (iii) For the corresponding quantum map, we try to replace as far as possible random “which path” decisions by a superposition of all available options. This is done in such a way that the absolute value square of the transition amplitudes agree with the corresponding classical transition probabilities. (iv) This must be done in such a way that Eq. (8) is fulfilled. As has been explained in [4] it is not an easy task to verify this condition. To lighten this task, in the next section we analyze the relationship that must exist between the Kraus operators of the quantum map $\Lambda_{\mathcal{Q}}$ and the matrix elements of the classical transition matrix $\mathcal{T}$.
@@ III Implementation in the Ising model
Consider a linear chain made up $N$ subsystems of two classical states, or spins. We assume the chain is embedded in a thermal bath at very high temperature. In this condition the chain is found in a disordered phase, i.e., each spin can be in one of its two possible classical states with equal probabilities. Then, a quench of temperature is applied from $T\rightarrow\infty$ to $T\rightarrow 0$, and the chain enters in a relaxation process to the equilibrium. This relaxation process can be modeled by the Glauber dynamics, that consist in the successive application of local stochastic maps, $\mathcal{T}^{(q)}$. The elements of the chain are labeled as: $s_{N},s_{N-1},\ldots,s_{0}$, and we denote the two states of each spin, $s_{q}\in\{0,1\}$, where 00 represents spin up and $1$ spin down. The local operation changes the state of the spin, $s_{q}\,(1\leq q\leq N)$, depending of the states of its near neighbors as follows
$\mathcal{T}^{(q)}:\quad s_{q}\longrightarrow\begin{cases}s_{q-1}&:\quad s_{q-1}=s_{q+1}\\ \begin{matrix}0\,(50\%)\\ 1\,(50\%)\end{matrix}&:\quad\mbox{otherwise}.\end{cases}$ (10)
The notation (50%) signifies that the state of the subsystem $s_{q}$ is assigned randomly. The elementary local operation can be expressed equivalently as a transition matrix in the subspace of the three spins, $s_{n-1},s_{n},s_{n+1}$, which basis is denoted as $\{|000\rangle,|001\rangle,|010\rangle,|011\rangle,|100\rangle,|101\rangle,|110\rangle,|111\rangle\}$,
@@
The goal of the quantum extension is to replace the statistical mixture in the configurations, $\{|001\rangle,|011\rangle\}$ and $\{|100\rangle,110\}$, by superposition of those configurations. Leading us to quantum evolutions close and not so close to the vicinity of the classical process. The single-spin local unitary operation that allows us to create these superpositions is the Hadamard gate
$H$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (12)
which transform the classical states $|0\rangle$ and $|1\rangle$ of a single spin as
$H|0\rangle$ $=\frac{(|0\rangle+|1\rangle)}{\sqrt{2}}=|+\rangle$
$H|1\rangle$ $=\frac{(|0\rangle-|1\rangle)}{\sqrt{2}}=|-\rangle$
[…]
$S$ $=\sigma_{x}H\equiv\frac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$ (13)
which action on the classical states is
$S|0\rangle$ =|=|$\rangle$= | + ⟩
[…]
In the matrix representation, $H$ and $S$ differ only in the position of a negative sign. This seemingly insignificant difference plays a very important role in the dynamics of the whole chain, as we will see below. This shows us that small local changes can generate important global effects when considering complex multipartite systems, to the extent that macroscopic observables can behave in unexpected ways.
To construct the quantum version(s) of the classical stochastic process, we follow the ideas presented in [4]. Thus, the local quantum operation, in the three-spin subspace, can be written as $\Lambda_{\mathcal{Q}}^{(q)}:=\{\mathcal{K}_{1},\mathcal{K}_{2}\}$, with
@@
The elements, $X_{j,k}$, of the embedded matrix $X$ of dimension $4\times 4$ are in general arbitrary complex numbers. In terms of $X$, the CPTP condition reduces to that the scalar product $\langle\vec{v}_{j}|\vec{v}_{k}\rangle=\delta_{jk}$ (for $j,k=1,\ldots,4$) where the vector $\vec{v}_{j}$ is defined by the $j$th column of $X$. In addition, $\Lambda_{\mathcal{Q}}^{(q)}$ is a quantum extension of $\mathcal{T}^{(q)}$ if the following relations are satisfied
$\left|X_{11}\right|^{2}+\left|X_{31}\right|^{2}$ $=1/2,$ $\left|X_{13}\right|^{2}+\left|X_{33}\right|^{2}$ $=1/2$
$\left|X_{21}\right|^{2}+\left|X_{41}\right|^{2}$ $=1/2,$ $\left|X_{23}\right|^{2}+\left|X_{43}\right|^{2}$ $=1/2$
$\left|X_{12}\right|^{2}+\left|X_{32}\right|^{2}$ $=1/2,$ $\left|X_{14}\right|^{2}+\left|X_{34}\right|^{2}$ $=1/2$
$\left|X_{22}\right|^{2}+\left|X_{42}\right|^{2}$ $=1/2,$ $\left|X_{24}\right|^{2}+\left|X_{44}\right|^{2}$ $=1/2\,.$
For the purpose of our analysis it is enough to consider all the elements of $X$ as reals. In Ref. [4] two basics quantum extensions have been defined denoted as $HAD-0$ and $SYH-0$, which in our construction can we identify as $X\in\{H_{0},S_{0}\}$ respectively, where
@@
$S_{0}$ $=\mathds{1}\otimes S=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-1&0&0\\ 1&1&0&0\\ 0&0&1&-1\\ 0&0&1&1\\ \end{pmatrix}\,.$ (16)
here $\mathds{1}$ stands for the identity operation for a single-spin. To approximate the classical process one can perform projective measurements in the standard basis with a certain rate, as shown in [4]. Here we follow a different approach, instead of projective measurements, we propose alternative versions of the quantum extension. At the beginning of the evolution all these quantum extensions are equivalent to each other, in the sense that they introduce approximately the same amount of coherence into the system, however, the behavior of the macroscopic observables for later times is very different, to the extent of being opposite with respect to the classical process. This apparent paradox is explained by observing different dissipation rates in the coherences of the different quantum extensions. This is a consequence of the local phases introduced by the operation $S$ along the chain during the evolution.
To create alternative quantum extensions to Eq.(15) and Eq.(16), we again rely on the elementary operations for a single spin, $H$ and $S$, but applied in different sub-spaces of the three-spin subsystem, as shown below: $X\in\{H_{1},H_{2},S_{2},S_{1}\}$ with the following definitions
$H_{1}$ $=\sigma_{x}\otimes H=\frac{1}{\sqrt{2}}\begin{pmatrix}0&0&1&1\\ 0&0&1&-1\\ 1&1&0&0\\ 1&-1&0&0\\ \end{pmatrix}\,,$ (17a)
$H_{2}$ $=H\otimes H=\frac{1}{2}\begin{pmatrix}1&1&1&1\\ 1&-1&1&-1\\ 1&1&-1&-1\\ 1&-1&-1&1\\ \end{pmatrix}\,,$ (17b)
$S_{2}$ $=S\otimes S=\frac{1}{2}\begin{pmatrix}1&-1&-1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ 1&1&1&1\\ \end{pmatrix}\,,$ (17c)
$S_{1}$ $=\sigma_{x}\otimes S=\frac{1}{\sqrt{2}}\begin{pmatrix}0&0&1&-1\\ 0&0&1&1\\ 1&-1&0&0\\ 1&1&0&0\\ \end{pmatrix}\,.$ (17d)

### v2

@@ Abstract
Recently, a novel construction scheme for generating quantum analogs of classical stochastic processes has been introduced. Here, we use this scheme in order to generate a large class of self-contained quantum extensions of a classical Markov chain process using symmetry operations. We show that the relaxation processes unfold very differently for the different quantum extensions. This is supported by monitoring the coherence, the probability of reaching the equilibrium, the decay of the number of domain walls and the purity. Unexpectedly, we find a rather ambiguous relation between the coherence measure based on the L1-norm and the speed of the relaxation process. Finally we find that the finite size scaling of the coherence measure exists for both short and long times and the value of the critical exponent is different for the short and long time.
@@ I Introduction
In this paper, we investigate the frontier between quantum and classical stochastic systems. On a fundamental level, stochasticity is not intrinsic to classical dynamics, it rather emerges due to the chaotic dynamics of many particle systems. In quantum mechanics, by contrast, genuine stochasticity appears as soon as measurement processes are included (wave function collapse). Although the measurement process can be avoided in many different ways, for instance by including the measurement apparatus into the quantum system to be described, at some point, it has to pop up, if meaningful experimental results are to be discussed [1, 2, 3].
@@
A gradual transition from quantum dynamics (deterministic and unitary) to (partially) stochastic processes is usually the subject of study in the theory of open quantum systems. In this case, information about the dynamics of the quantum system is leaking out into the environment, a process which may be seen as applying measurements continuously with low probabilities. The characterization of such Markovian processes has been established in Refs. [4, 5, 6], and is by now, well established [7] and used in countless applications.
@@
Recently, an alternative approach (to describe systems intermediate between quantum unitary and classical stochastic) has been proposed [8]. There, one starts from the classical stochastic process, and asks for the (partially) quantum processes (quantum extensions) in its vicinity. This approach leads to processes which are close to classical but still quantum. In other words, they operate with finite, but very low amounts of coherence, as measured by the L1-norm based measure introduced in [9]). Surprisingly, however, these processes show macroscopic behavior that is different (even the scaling with system size) as compared to the classical mother process.
@@
The finding of stochastic processes which are close to classical (in terms of the coherence measure) and still show very different properties on the macroscopic scale (thermodynamic limit) is important for many reasons : (i) biochemical processes in warm and hostile environments are capable of producing sufficient coherence to make transport processes efficient (light harvesting, etc.) [10, 11] (ii) Superior quantum computation with low-coherence quantum computers, to name two [12].
@@
In the present contribution, we construct a large self-contained set of quantum extensions of the classical relaxation process in the Ising model. These extensions are generated from two single qubit operations, the Hadamard and the NOT gate. We then investigate the resulting processes by measuring several quantities as a function of time: (i) the probability to find the system in the ground state; (ii) the decay of the number of domains; (iii) the purity; and (iv) the coherence. We find that the amount of coherence is related to the degree to which the quantum extension deviates from the purely classical process, and we show the finite size scaling for its decay.
Section II, briefly introduces the concept of classical Markov processes. Subsequently, the concept of a quantum Markovian process and its connection with the formalism of open quantum systems is introduced. Finally, a brief explanation of the concept of quantum extension is given, which allows us to study quantum versions in the vicinity of a classical process. In Section III, the method of quantum extensions is applied to an Ising chain subject to a zero-temperature quench, and a strategy for constructing multiple quantum extensions is shown. In the first part of section IV, exact numerical simulations are performed for chains of length $N=12$, and the behavior of several macroscopic observables is analyzed. Subsequently, the scaling behavior of coherence is investigated for spin chains up to length $N=20$. In Section V, conclusions are presented and some perspectives have been discussed.
@@ II Classical and quantum stochastic processes
Following [8], we consider classical stochastic processes with a finite sample space, $\{|j\rangle\}_{j=1,\ldots,d}$, where each $|j\rangle$ represents a possible configuration of the physical system. In this case, the dynamics of the system can be characterized by a sequence of stochastic transition matrices, $\{\mathcal{T}(n)\,|\,n\in\mathbb{N}_{0}\}$, such that $\bm{p}(n+1)=\mathcal{T}(n)\;\bm{p}(n)$, i.e.
 $\forall\;1\leq i\leq d\quad:\quad p_{i}(n+1)=\sum_{j=1}^{d}\mathcal{T}_{ij}(n)\;p_{j}(n)\;.$ (1)
Here, the component, $p_{j}(n)$, of the vector $\bm{p}$ denotes the probability that the system is found in the configuration $|j\rangle$ at the discrete time $n$. The matrix element $\mathcal{T}_{ij}(n)$ is the conditional probability that the systems will be found in the configuration $|i\rangle$ at time $n+1$, provided it is in configuration $|j\rangle$ at time $n$. That means that the column vectors of any transition matrix have non-negative entries which sum up to one. This guaranties the conservation of probability along the process.
The stochastic process can be generated from the composition of subsequent transition matrices, realized by matrix multiplication [13]. That is, with $n\geq m\geq 0$:
$\bm{p}(n+1)=\mathcal{T}(n)\;\mathcal{T}(n-1)\;\cdots\;\mathcal{T}(m)\;\bm{p}(m)\;.$ (2)
Such a process is also called a “Markov chain”.
@@ II.1 Quantum processes
In quantum mechanics, the analog of a Markov chain is conveniently described within the framework of quantum channels [7, 14, 15]. Here, the probability vectors are replaced by density matrices $\varrho\in\mathcal{S}(\mathcal{H})$ [16, 17], where $\mathcal{H}$ is the Hilbert space of all linear combinations of classical configurations $\{|j\rangle\}_{j=1,\ldots,d}$, equipped with the scalar product where $\langle i|j\rangle=\delta_{ij}$.
The space $\mathcal{S}(\mathcal{H})$ is made up of all convex combinations of one-dimensional projectors in $\mathcal{H}$. That is
$\mathcal{S}(\mathcal{H})=\Big{\{}\varrho={\textstyle\sum_{j}}\,p_{j}\;\frac{|\psi_{j}\rangle\langle\psi_{j}|}{\|\psi_{j}\|^{2}}\;\Big{|}\;{\textstyle\sum_{j}}\,p_{j}=1\,,\;\psi_{j}\in\mathcal{H}\,\Big{\}}\;.$ (3)
Equivalently, we may say that $\varrho\in\mathcal{S}(\mathcal{H})$ if and only if $\varrho$ is a positive, Hermitian operator on $\mathcal{H}$, with unit trace. A quantum Markov chain can then be constructed from quantum channels, i.e. completely positive and trace preserving (CPTP) linear maps on $\mathcal{S}(\mathcal{H})$ [18]. Hence, instead of evolving probability vectors, we are now evolving density matrices in time, such that
$\varrho(n+1)$ $=\Lambda_{\mathcal{Q}}(n)\circ\cdots\circ\Lambda_{\mathcal{Q}}(m)\,[\varrho(m)]\;,$ (4)
where $\circ$ denotes composition [14, 15] and the discrete times are ordered as $n\leq m\leq 0$.
In what follows, we will make use of the fact that every CPTP map admits a sum representation in terms of Kraus operators [19]. Then, for each element in the Markov chain given in Eq.(4), we can write
$\Lambda_{\mathcal{Q}}(n)[\varrho(n)]$ $=\sum_{\alpha}\mathcal{K}_{\alpha}(n)\,\varrho(n)\,\mathcal{K}_{\alpha}^{\dagger}(n)\,,$ (5)
where the Kraus operators, $\{\mathcal{K}_{\alpha}(n)\}$, satisfy the condition
$\forall\;n\in\mathbb{N}_{0}\quad:\quad\sum_{\alpha}\mathcal{K}_{\alpha}^{\dagger}(n)\mathcal{K}_{\alpha}(n)=\mathds{1}\;.$ (6)
Note that the quantum Markov chains defined here, include classical processes such as the ones defined in Eq.(2). To see this, we first express the probability vector, $\bm{p}$, as a diagonal density matrix, $\bar{\varrho}$, and subsequently replace the transition matrices, $\mathcal{T}(n)$, by the quantum channels $\Lambda_{\mathcal{T}}(n)$,
$\bar{\varrho}(n+1)=\Lambda_{\mathcal{T}}(n)[\bar{\varrho}(n)]=\sum_{i,j}\mathcal{K}_{ij}(n)\bar{\varrho}(n)\mathcal{K}_{ij}^{\dagger}(n),$ (7)
with the following definition for the Kraus operators
$\Big{\{}\mathcal{K}_{ij}(n)$ $=\sqrt{\mathcal{T}_{ij}(n)}{|i\rangle\langle j|}\Big{\}}_{1\leq i,j\leq d}\,,$ (8)
where it is straight forward to show that condition (6) is always satisfied. Note that the maps $\Lambda_{\mathcal{T}}(n)$ belong to the set of incoherent CPTP maps, which are neither been able to create nor detect coherence [9, 20].
@@ II.2 Quantum extensions
We are interested in quantum processes which reduce to a given classical process when observed or measured sufficiently often. We name such processes “quantum extensions” of the given classical stochastic process. The idea is that if we observe a classical stochastic process, we may think that its stochastic nature really comes from a quantum process, which has lost its coherence due to direct or indirect (i.e. coupling to some environment) measurements. The quantum extensions we are interested in, may then be seen as those quantum processes, where measurements are not sufficiently complete or frequent, such that the resulting process is only partially incoherent. This idea can be clarified by the following formal definition [8] :
Definition:
@@
where $\mathcal{P}$ denotes a complete measurement of the set of classical configurations (i.e. , the basis states $|j\rangle$).
To find quantum extensions for a given classical Markov chain, we use the following guiding principles. (i) We try to find quantum extensions for each stochastic map $\Lambda_{\mathcal{T}}(n)$, individually. (ii) For the corresponding quantum maps, we try to replace as far as possible random “which path” decisions, encoded in the matrix elements of $\mathcal{T}_{ij}(n)$, by superpositions of all available options. (iii) In order to obtain a valid quantum extension, we make sure that Eq. (9) is fulfilled.
As explained in [8] it is not an easy task to verify this condition. To shed some light on this task, in the next section we analyze the relationship that must exist between the Kraus operators of the quantum map $\Lambda_{\mathcal{Q}}$ and the matrix elements of the classical transition matrix $\mathcal{T}$.
@@ III Implementation in the Ising model
Consider a linear chain made up $N$ subsystems of two classical states, or spins. We assume the chain is embedded in equilibrium with a thermal bath at very high temperature ($T\rightarrow\infty$ ). In this condition, the chain is found in a disordered phase, i.e., each spin can be in one of its two possible classical states with equal probabilities. Then the system is quenched to zero temperature ($T=0$) and as a result the chain enters a relaxation process to reach equilibrium. We model this relaxation process using the Glauber dynamics [21] that have been widely used to study the zero-temperature dynamics for classical spin systems [22, 23, 24]. The process consists of the successive application a global stochastic map $\mathcal{T}(n)=\mathcal{T}_{\rm all}$, independent of $n$, which is a uniform mixture of local stochastic maps, $\mathcal{T}^{(q)}$, to be explained below. In other words,
$\mathcal{T}_{\rm all}=\frac{1}{N}\sum_{q=1}^{N}\mathcal{T}^{(q)}\;.$ (10)
This construction describes the procedure where one selects at random one of the spins in the chain, and then updates the spin, according to Glauber’s acceptance criterion. In this scenario, one Monte Carlo time step (MCS) corresponds to $N$ applications of $\mathcal{T}_{\rm all}$, which in turn corresponds to $N$ local updates (including random selection) in the Glauber algorithm.
To define the local update operation, we enumerate the spins from $q=1$ to $N$, denote a classical configuration of the spin chain by $|\vec{s}\rangle=|s_{1},s_{2},\ldots,s_{N}\rangle$ and adopt periodic boundary condition, such that $s_{0}=s_{N}$ and $s_{N+1}=s_{1}$. The two possible states of each spin are $s_{j}\in\{0,1\}$, where 0 (1) denotes the spin pointing upward (downward). The local operation $\mathcal{T}^{(q)}$ involves the spin $q$ and its immediate neighbors. In the truncated configurational basis $\{|s_{q-1},s_{q},s_{q+1}\rangle=|000\rangle,|001\rangle,|010\rangle,|011\rangle,|100\rangle,|101\rangle,|110\rangle,|111\rangle\}$, it reads
@@
To construct the quantum version(s) of the classical stochastic process, we translate Eq. (10) into the quantum channel setting, and search for quantum extensions of the local update operations $\mathcal{T}^{(q)}$. That is,
$\Lambda_{\mathcal{Q}}^{\rm all}=\frac{1}{N}\sum_{q=1}^{N}\Lambda_{\mathcal{Q}}^{(q)}\;,$ (12)
where the channels $\Lambda_{\mathcal{Q}}^{(q)}$ should be constructed in such a way that random choices applied in case of the configurations, $\{|001\rangle,|011\rangle\}$ and $\{|100\rangle,110\}$, are replaced by their respective superpositions. To this end, we assume that $\Lambda_{\mathcal{Q}}^{(q)}$ can be described by a small number of Kraus operators. Then we investigate whether Eq. (6) (completeness of the Krauss decomposition) and Eq. (9) (validity as a quantum extension) can be fulfilled.
In the case of only one Kraus operator, that operator must be a unitary. In the present case, this is not possible because the classical map $\mathcal{T}_{q}$ maps two configurations $|000\rangle$ and $|010\rangle$ onto the same output configuration, $|000\rangle$. We will therefore concentrate on the case of two Kraus operators.
$\Lambda_{\mathcal{Q}}^{(q)}[\varrho]=\mathcal{K}_{1}\,\varrho\,\mathcal{K}_{1}^{\dagger}+\mathcal{K}_{2}\,\varrho\,\mathcal{K}_{2}^{\dagger}\;.$ (13)
In that case, the most general form of the two operators is as follows:
@@
Here we collect the potentially free parameters into a $4$$\times$$4$ matrix $X$. We then find that in order to fulfill Eq. (6) the column vectors of $X$ must be orto-normal, and in order to fulfill Eq. (9), it must hold
[…]
For the purpose of our analysis it is enough to consider all the elements of $X$ as reals.
Generating set of quantum extensions:
As explained at the previous section, the core of the quantum construction is based on two elementary operations. $\sigma_{x}$ that flip spins and the Hadamard gate $H$ that replace the statistical mixtures by superpositions. One can observe that combinations of these two operations, $\sigma_{x}H$ or $H\sigma_{x}$, are also valid that creates superpositions. Then, one wonders if it is possible to construct different quantum extensions based in these elemental operations. Considering the minimal set $g:=\bm{\{}\sigma_{x},H\bm{\}}$ along with the multiplication operation, it is not difficult to show that the following group can be constructed
$\mathcal{G}:=$±$\bm{\{}\pm\mathds{1},\pm\sigma_{x},\pm i\sigma_{y},\pm\sigma_{z},\pm H,\pm H\sigma_{x},\pm\sigma_{x}H,\pm\sigma_{x}H\sigma_{x}\bm{\}}\,.$ (16)
You should notice that $\mathcal{G}$ contains the identity, $\mathds{1}$, and given that every element of the set is unitary its inverse exists. To generate different quantum extensions in terms of the matrix $X$, here we propose that $X$ is an element of the set $\mathcal{G}^{\otimes 2}$, whenever the conditions to be a quantum extension defined in the previous section are satisfied. The results are summarized in Table 1. Take for example, the two basic quantum extensions defined in Ref. [8], HAD-0 and SYH-0 which in our construction can be identified as $X\in\{H_{0},S_{0}\}$ respectively, where
@@
$S_{0}$ $=\mathds{1}\otimes\sigma_{x}H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-1&0&0\\ 1&1&0&0\\ 0&0&1&-1\\ 0&0&1&1\\ \end{pmatrix}\,.$ (18)
Table 1: Quantum extensions generated by the combination of the positive elements in $\mathcal{G}$. The check-mark stands for the valid quantum extensions.
$\bigotimes$ $\mathds{1}$ $\sigma_{x}$ $i\sigma_{y}$ $\sigma_{z}$ $H$ $H\sigma_{x}$ $\sigma_{x}H$ $\sigma_{x}H\sigma_{x}$
$\mathds{1}$
$\sigma_{x}$
$i\sigma_{y}$
$\sigma_{z}$
$H$
$H\sigma_{x}$
$\sigma_{x}H$
$\sigma_{x}H\sigma_{x}$
Equivalent quantum extensions:
In the Table 1, we observe all the possible combinations that produces valid quantum extensions, but several of them are equivalent, such that we can reduce the set from 32 to the following set of 12 different quantum extensions
$\bm{\Big{\{}}\mathds{1}$ $\otimes H,$ $\mathds{1}$ $\otimes\sigma_{x}H$
$\sigma_{x}$ $\otimes H,$ $\sigma_{x}$ $\otimes\sigma_{x}H$
$i\sigma_{y}$ $\otimes H,$ $i\sigma_{y}$ $\otimes\sigma_{x}H$
$\sigma_{z}$ $\otimes H,$ $\sigma_{z}$ $\otimes\sigma_{x}H$
$H$ $\otimes H,$ $H$ $\otimes\sigma_{x}H$
$\sigma_{x}H$ $\otimes H,$ $\sigma_{x}H$ $\otimes\sigma_{x}H\bm{\Big{\}}}$ (19)
At the beginning of the evolution all these quantum extensions introduce approximately the same amount of coherence into the system, however, the behavior of the macroscopic observables for later times is very different, to the extent of being opposite with respect to the classical process. This apparent paradox is addressed by observing different dissipation rates in the coherences of the different quantum extensions.